OpenStudy (anonymous):

John is currently 26 years older than Mary. 12 years from now , John will be twice as old as Mary. What is John's currently age?

OpenStudy (anonymous):

let age of mary=x age of john=y John is currently 26 years older than Mary y=x+26 ----(A) 12 years from now , John will be twice as old as Mary y+12=2x -----(B) from (A) and (B), y=x+26 and y=2x-12 so x+26=2x-12 so x=38 and y=x+26=38+26=64 x=38, y=64